Circular No. 7434

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000ck IN IC 4355 S. Benetti, R. Cosentino, J. Licandro, F. Paulli, M. Pedani, G. Trancho,

A. Zacchei, and E. Giro, Telescopio Nazionale Galileo (TNG); P. Conconi, G. Crimi, and E. Molinari, Osservatorio Astronomico di Brera; and A. Caproni, M. Comari, C. Corte, S. Monai, and M. Pucillo, Osservatorio Astronomico di Trieste, report: “A fully reduced CCD spectrum (range 333.3–722.3 nm, resolution 1 nm), obtained on May 31.03 UT with the TNG reflector (+ D.o.lo.res spectrograph), tentatively shows SN 2000ck (cf. IAUC 7431) to be a type-II supernova near maximum light. The spectrum is dominated by a very blue continuum (T_{bb} ∼ 17 000 K), upon which we identify (even if faint) Hδ, Hγ, and Hβ lines with relatively narrow P-Cyg profiles (from which we derive an expansion velocity of ∼ 6200 km/s). There is no sign of a broad Hα feature. From weak interstellar lines due to the Na I D absorption feature (EW ∼ 0.12 nm) seen at ∼ 605.3 nm, we derive a parent-galaxy recession velocity of ∼ 8150 km/s.”

S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics; and P. Garnavich, University of Notre Dame, write: “Spectra of SN 2000ck, taken by P. Berlind and Garnavich on May 27.4, 28.3, 29.2, 30.2, and 31.3 UT with the Whipple Observatory 1.5-m telescope (+ FAST spectrograph), show it to be an unusual type-II supernova at an early epoch. The spectra exhibit a blue continuum with narrow emission lines of [O II], Hγ, Hβ, [O III], [O I], [N II], Hα, and [S II] from a superimposed H II region at the recession velocity of IC 4355 (8048 km/s in the NASA/IPAC Extragalactic Database). From the supernova itself, there are broad but very weak absorption features of Hβ (expansion velocity 6700 km/s) and He I (rest 587.6 nm, expansion velocity 5300 km/s) that have developed only very slowly. Strong interstellar Na I absorption (with equivalent width 0.1 nm) at the host-galaxy redshift implies significant extinction by dust.”

A. V. Filippenko and A. C. Coil, University of California at Berkeley, communicate: “A CCD spectrum (range 320–1000 nm) of SN 2000ck obtained on May 31 UT with the Shane 3-m reflector at Lick Observatory exhibits a blue, nearly featureless continuum. It resembles the spectrum of the peculiar type-II supernova 1993J at very early times (Filippenko 1997, ARAA 35, 309), but the type-II classification is not yet certain.”

COMET C/1999 S4 (LINEAR) Visual m_{1} estimates by N. Biver, Oahu, HI (0.26-m reflector): May

25.61 UT, 11.3; 29.61, 11.0.

2000 June 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green